OpenStudy (anonymous):

4x - 2y = 5 3x + 3y = 7 What is the system determinant? -18 -6 6 18

OpenStudy (callisto):

|4 -2 | |3 3 | = 4(3) - (-2)(3) = 12 + 6 =?

OpenStudy (anonymous):

18